You have survey data from a sample of 20 people. What t critical statistic would you use for a left-tailed test with an alpha of .01?

Round to the third decimal place.
Answer #1

solution

it is left tailed test

n =20

Degrees of freedom = df = n - 1 =20 - 1 = 19

= 0.01

t ,df = t0.01,19 = 2.539 ( using student t table)